AN IMPROVED CHIRPED PULSE FTMW ANALYSIS OF THE STRUCTURES OF PHENOL DIMER AND TRIMER

\underline{NATHAN A. SEIFERT}, CRISTÓBAL PÉREZ, AMANDA L. STEBER, DANIEL P. ZALESKI, JUSTIN L. NEILL, BROOKS H. PATE, Department of Chemistry, University of Virginia, Mc-Cormick Rd., Charlottesville, VA 22904-4319; ALBERTO LESARRI, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47001 Valladolid, Spain.

With the recent improvements for chirped pulse FTMW (CP-FTMW) spectroscopy between 2-18 GHz, substitution structures of molecules and clusters with more than 10 heavy atoms are becoming routine. While previous CP-FTMW results for phenol dimer reported at this conference by Steber et al.^{a} necessitated reduced-band measurements in order to achieve the sensitivity to detect the carbon isotopologues, the latest improvements for the 2-8 GHz arrangement have enabled full band detection of all 12 ^{13}C and 2 ^{18}O isotopologues of phenol dimer in natural abundance, with improved fits for all detected species. In addition, the added sensitivity of this new 2-8 GHz configuration has enabled a full carbon substitution structure of phenol trimer. The experimental structure of phenol trimer, in agreement with the M06-2X/6-311++g(d,p) ab initio structure, is a C_{3} oblate symmetric top with 21 heavy atoms; however, all possible isotopic substitutions are off-symmetry axis, so the resulting detected isotopologues have been fit as c-type prolate asymmetric tops. Use of Kraitchman′s equations for structural determination of a symmetric top molecule require some assumptions from the ab initio structure for the complete r_{s} structure of the trimer. A detailed summary of these methods, as well as the microwave results for both species, will be presented.

aA. L. Steber, J. L. Neill, D. P. Zaleski, B. H. Pate, A. Lesarri. 67th OSU Int. Symp. On Mol. Spectrosc., Columbus, OH, 2012, MH13.